# Building Better: Avoiding Pitfalls in Developing Language Resources when Data is Scarce

### v5

@@ Abstract
Language is a symbolic capital that affects people’s lives in many ways Bourdieu (1977, 1991). It is a powerful communication means that accounts for identities, cultures, traditions, and societies in general. Hence, data in a given language should be viewed as more than a collection of tokens. Good data collection and labeling practices are key to building more human-centered and socially aware technologies. While there has been a rising interest in under-resourced languages within the NLP community, work in this space has to overcome unique challenges such as data scarcity and access to suitable annotators. In this paper, we collect feedback from those directly involved in and impacted by NLP artefacts for medium and low-resource languages. We conduct a quantitative and qualitative analysis of the responses and highlight the main issues related to (1) data quality such as linguistic and cultural data suitability; and (2) the ethics of common annotation practices such as the misuse of online community services. Based on these findings, we make several recommendations for the creation of high-quality language artefacts that reflect the cultural milieu of its speakers, while simultaneously respecting the dignity and labor of data workers.
@@ 1 Introduction
Figure 1: The main stakeholders and their role in a data annotation project. We reached out to the CL community in our survey. It is designed for NLP researchers and practitioners who have worked on low- to medium-resource languages. Some of the questions focus on the perspectives of the subset highlighted in the figure, i.e., speakers who focus on their own languages.
Over the past few years, there has been an increasing interest in improving the scope of NLP research with more human-centered design choices Kotnis et al. (2022) and social awareness Yang et al. (2024). As language technologies depend on the quality of the data Hirschberg and Manning (2015) and their alignment with the needs of the speakers, researchers, and other users, the perspectives of these different stakeholders are key to high-quality tools and resources. That is, data selection, collection, annotation, and design choices are traditionally made by the researchers who develop the different artefacts. However, the involvement of those whose native languages are in question is paramount to better design practices Bird and Yibarbuk (2024) since language is part of their culture and identity Bourdieu (1991). That said, when dealing with low to mid-resource languages in NLP, researchers often make use of the datasets available without necessarily looking into their quality or questioning their suitability, mainly due to resource scarcity. Although progress in NLP for English and other high-resource languages has led to improving standards for corpora quality control and research practices Gebru et al. (2021); Bender and Friedman (2018); Mohammad (2022), one cannot claim the same about the data sources and prevailing practices for low- to mid-resource languages given the current research scope in the field Joshi et al. (2020). Therefore, the NLP artefacts developed for under-resourced languages and underrepresented cultures often suffer from a lack of social considerations and cause over-generalisations due to the over-reliance on data and tools that fail to incorporate the predominant linguistic and cultural features of a given language Bender and Friedman (2018), which may hinder critical progress. This can further lead to inequality Blasi et al. (2022); Held et al. (2023), sub-optimal experiences with language technologies, and could reinforce a legacy of language hierarchy Kahane (1986).
In this position paper, we shed light on the current limitations of NLP research for low- to mid-resource languages in terms of appropriate data collection, ethical annotation practices, and overall data quality. We reached out to the NLP community involved in NLP projects on under-served languages and conducted a survey to report on the common incentives, limitations, applied norms, and practices (see Figure 1). We outline the survey and present its results. Finally, based on the survey responses, we provide a set of recommendations that focus on (1) fairness and centering of the speakers of the language, (2) choosing suitable data sources, (3) setting fair and realistic expectations when recruiting annotators, and (4) avoiding cultural misrepresentation.
@@ 2 Related Work
Work on ethical practices in AI, ML, and NLP research covers a variety of topics, such as artefact documentation Bender (2011); Bender and Friedman (2018); Gebru et al. (2021); Rogers et al. (2021); Mohammad (2022) and recommendations for best practices Hollenstein et al. (2020); Mohammad (2023). Those that focus on low-resource languages are centered on the general state of NLP research in the area Held et al. (2023); Joshi et al. (2020); Blasi et al. (2022); Doğruöz and Sitaram (2022), limitations in specific tasks such as machine translation Mager et al. (2023), LLM research Mihalcea et al. (2024), or on the essential question of including people whose languages are in question Mager et al. (2023); Bird (2020, 2022); Bird and Yibarbuk (2024); Lent et al. (2022). Such work sheds light on the peculiarities of low-resource languages with the majority being vernacular languages rather than institutionalised or written Bird and Yibarbuk (2024); Bird (2024). They further advocate for language communities to take over their languages Schwartz (2022); Markl et al. (2024); Mihalcea et al. (2024). For instance, Bird and Yibarbuk (2024) focus on how experts (e.g., linguists, computer scientists) interact with the language communities using participatory design approaches Winschiers-Theophilus et al. (2010), and Cooper et al. (2024) provide recommendations on how to engage with indigenous communities without merely focusing on accuracy. Doğruöz and Sitaram (2022) further point out the importance of not treating language technologies for low-resource languages as scaled-down versions of high-resource ones and account for their specificity. Similarly, Adebara and Abdul-Mageed (2022) make related claims with a focus on features that are peculiar to African languages (e.g., tones). In addition to the language speakers, other work focuses on users such as Blaschke et al. (2024) who highlight the needs of dialect speakers and the importance of involving them in designing language technologies. Moreover, Yang et al. (2024) define social awareness and advocate for refraining from treating language in NLP as a computational problem only. In this paper, we strengthen the above discussion by shifting the focus to the practical challenges faced by NLP researchers and practitioners working on low and mid-resource languages by borrowing practices from social science Cetina (1999) to study the methodological practices and issues in the field. To the best of our knowledge, there is limited work investigating NLP research for low-resource languages while trying to connect to online communities, except for three case studies of participatory frameworks (e.g., Masakhane), and work by Lent et al. (2022), who analyse 38 responses by Creole speakers on their experiences with language technologies. By analysing the respondents’ feedback, we present practical recommendations that emphasise transparency and ethically grounded practices for building more human-centered NLP artefacts for under-served languages.
@@ 3 Survey
Our main goal is to investigate the current issues and problematic practices in NLP research for low- to mid-resource languages and provide potential solutions. Therefore, we reached out to the NLP community (i.e., *CL networks) from June to October 2024 on X, LinkedIn, Google groups, and Slack channels of online NLP communities, and by direct emails. We targeted researchers and annotators working on low- and mid-resource languages, language variants, dialects, and vernaculars, and surveyed how research is conducted. Participants report on common practices, incentives, and issues that stand out. Then, we present a quantitative and qualitative analysis of the responses.
3.1 Respondents
Respondents are NLP researchers and practitioners involved in the data collection, annotation, model construction, or other research questions related to under-served languages.
3.2 Survey Structure
We ask the respondents about (1) their previous experiences in the area, (2) current problems and limitations related to their language(s) of interest, (3) the motivation behind their involvement in various projects, and (4) how they were credited for tasks that are often specific to low-resource languages, e.g., compensation for annotations done via online community forums or participatory frameworks.
Note that we left it to the respondents to decide on what represents a mid- to low-resource language since there is no standard definition. E.g., most researchers would agree that Tamasheq is a low-resource language, while they may not have the same opinion about Malaysian.
3.2.1 General Questions
Respondents could optionally fill in their names and contact information for a potential follow-up. Then, they were asked about:
  • the language(s) they work on,
  • the project(s) they were involved in,
  • whether they are/were part of any online community (i.e., participatory research frameworks),
  • whether the project(s) they worked on are from industry, academia, or both,
  • the kind of NLP tools that are or would be relevant and useful in their language(s) of interest,
  • the reason(s) why they work on this/these language(s).
3.2.2 Reporting on Incentives and Potential Limitations
We investigate the common reasons why researchers work on low-resource languages. Therefore, we ask the participants to report on:
  • the incentive(s) for working on their language(s) of interest,
  • the incentive(s) for working on specific projects.
As we are aware of potential drawbacks in NLP for mid- to low-resource languages Blasi et al. (2022), we examined whether the respondents work in the area due to any limitations observed in available NLP tools in their language(s) of interest. Note that these questions were optional as researchers may work on low- to mid-resource languages for various other reasons. We asked the participants to report on:
3.2.3 Reporting on Credit Attribution
We asked the respondents about how often they were properly credited for their work. Further, as reaching out to online participatory frameworks is common to projects that focus on under-resourced languages, we asked whether the participants were involved in past projects through participatory frameworks. This is because involving communities in NLP and ML projects is relatively new to the field and can therefore be abused (purposefully or not) due to the lack of clear standards regarding data workers in such contexts. Therefore, our questions were the following:
4 Findings
@@
We received 81 responses from researchers working on a wide range of languages and language families. Even though including contact information was optional, more than 90% of the respondents chose not to reply anonymously, and 80% asked for updates on the project. Table 1 shows the distribution of responses to questions on project affiliations, the tasks in which the respondents were involved, and their motivations for working on under-resourced languages. Note that percentages do not sum up to one as respondents could report on more than one project. That is, participants could be involved in several tasks and projects. As shown in Table 1, most participants were involved in dataset curation mainly motivated by scientific interest or curiosity, and for building language technologies because of observed limitations in resources dedicated to their language(s) of interest.
@@ 4.1.1 Projects
The respondents could report on one or many projects they have been involved in. As shown in Table 1, Most respondents have worked on academic projects, with a third on collaborations between industry and academia or both types of projects. [^1] Note that although ¿50% of the respondents named the projects they participated in, we do not disclose it to protect the anonymity of our respondents.
@@ 4.1.2 Languages
Among the 81 responses, the respondents worked on >70 low-resource languages they specifically named (see Appendix). Figure 2 illustrates the main locations where these languages are spoken. These include variants, dialects, and vernaculars (e.g., country-specific Arabic dialects), truly low-resource languages (e.g., Welsh, Yoreme Nokki, Setswana), and low- to medium-resource ones (e.g., Amharic, Indonesian). In addition, about 12% of the respondents reported working on language families and language branches such as South Asian languages, all Gaeilige dialects, or Arabic/English variations. A high percentage of the respondents work on high-resource languages in parallel.
@@ 4.2 Incentives and Potential Limitations
When asking the respondents about why they work on NLP for under-resourced languages, we provide them with a checklist from which they could choose more than one option or add their own entry. We report on the frequent motivations and practices that are only adopted in low- to mid-resource settings due to, e.g., data scarcity. We identify problematic instances and analyse the possible reasons behind some.
@@
When further examining the common motivations, we report more detailed numbers in Figure 3. Among those who were motivated by scientific curiosity or interest in Table 1 there were those whose interest was in NLP/CL/ML research (68%) and those whose interest was in languages (68%). Note that the two are not mutually exclusive.
Moreover, for the respondents whose motivation was building language technologies, most of them were more interested in building technologies for their own language(s) (60%) as opposed to building technologies for as many languages as possible (38%). This is particularly interesting as it constitutes evidence of the power of language as a symbolic capital Bourdieu (1991), which can sometimes manifest in the feeling of “a duty” that one has towards their language. Other frequent motivations include marked limitations in language resources and tools in the language(s) of interest (60%) and the willingness to contribute to LLM research (59%).
@@ 4.2.2 Qualitative Analysis of the Limitations
Respondents on getting credit for projects they were involved in. We provided the respondents with free text sections where they could report examples of tools or resources that suffer from the limitations that they mentioned to justify their choices. When manually processing the answers, we noticed the following recurring topics:
@@
    Limitations related to the currently available resources: such as their unavailability, small size, limited representativeness, and quality.
@@
    Fundamental problems related to NLP research on low- to mid-resource languages: such as the lack of funding often due to the “low prestige” language dilemma—the false notion that some languages or language varieties are more important than others. Interestingly, this was equally observed in projects from academia and industry.
@@
Lack of representativeness and unnaturalness of the data were commonly reported in the responses. The respondents reported a lack of awareness of language variants and cultural aspects when building a language-specific artefact; the reliance on the standardised version of a given language due to power dynamics (more power in the hands of well-funded institutions and established researchers); the presence of offensive utterances in the data due to a lack of data filtering; and potentially wrong assumptions about a language or a culture. For instance, the time-specific context and usage of some languages, such as ancestral ones (e.g., Coptic), have considerably changed and one has to take these facts into account. In addition, datasets may be collected from inadequate sources or could be aligned with Western values, standards, or expectations. This can be due to power differentials or a lack of deeper examination carried along with locals and native speakers. Finally, researchers rely on personal connections as it is hard to impossible to find fluent speakers of low- to mid-resource languages on commonly used annotation platforms such as Amazon Mechanical Turk and Polific. Added to this reason, the lack of funding leads researchers to turn to participatory frameworks. This practice has been at the center of major NLP contributions in recent years Birhane et al. (2022). However, despite its benefits for people with common research interests, the absence of well-established standards puts vulnerable community members at risk as their efforts may not be properly recognised.
@@
Many respondents reported that conducting research in low- to mid-resource languages often entails high costs of data curation and potential outreach to local communities. Further, when resources for an under-resourced language exist, they are often not freely available.

### v6

@@ Abstract
Language is a form of symbolic capital that affects people’s lives in many ways Bourdieu (1977, 1991). As a powerful means of communication, it reflects identities, cultures, traditions, and societies more broadly. Therefore, data in a given language should be regarded as more than just a collection of tokens. Rigorous data collection and labeling practices are essential for developing more human-centered and socially aware technologies. Although there has been growing interest in under-resourced languages within the NLP community, work in this area faces unique challenges, such as data scarcity and limited access to qualified annotators.
In this paper, we collect feedback from individuals directly involved in and impacted by NLP artefacts for medium- and low-resource languages. We conduct both quantitative and qualitative analyses of their responses and highlight key issues related to: (1) data quality, including linguistic and cultural appropriateness; and (2) the ethics of common annotation practices, such as the misuse of participatory research. Based on these findings, we make several recommendations for creating high-quality language artefacts that reflect the cultural milieu of their speakers, while also respecting the dignity and labor of data workers.
@@ 1 Introduction
Figure 1: The main stakeholders and their roles in a data annotation project: We conducted a survey and reached out to the CL community, specifically NLP researchers and practitioners who have worked on medium- to low-resource languages. Some of the questions focus on the perspectives of a subset of stakeholders highlighted in the figure—namely, speakers who focus on their own languages.
Over the past few years, there has been growing interest in making NLP research more human-centered Kotnis et al. (2022) and socially aware Yang et al. (2024). As language technologies are deeply dependent on data quality Hirschberg and Manning (2015) and its alignment with the needs of speakers, researchers, and other users, incorporating diverse stakeholder perspectives is essential for building high-quality tools and resources. Traditionally, decisions around data selection, collection, annotation, and model design have been made primarily by researchers. However, the involvement of native speakers, whose languages are at the core of these artefacts, is paramount to better design practices Bird and Yibarbuk (2024), since language is inseparable from culture and identity Bourdieu (1991). Despite this, when it comes to mid- to low-resource languages, researchers often resort to using any available datasets, seldom scrutinising their quality or relevance due to resource scarcity. While NLP for English and other high-resource languages has benefited from improved standards for corpus quality and ethical research practices Gebru et al. (2021); Bender and Friedman (2018); Mohammad (2022), these standards are not consistently extended to under-resourced languages Joshi et al. (2020). Therefore, NLP artefacts for underrepresented languages often lack cultural and linguistic grounding, leading to false generalisations and systemic shortcomings Bender and Friedman (2018). This disconnect may hinder meaningful progress and perpetuate inequalities Blasi et al. (2022); Held et al. (2023), produce sub-optimal user experiences, and reinforce long-standing language hierarchies Kahane (1986).
In this position paper, we highlight the current limitations of NLP research for mid- to low-resource languages, specifically in terms of data collection, ethical annotation practices, and overall data quality. We reached out to the CL community involved in NLP projects on under-served languages and conducted a survey to report on the common incentives, limitations, applied norms, and practices (see Figure 1). We present the survey’s results and provide a set of recommendations based on the responses, focusing on (1) fairness and centering of the language speakers, (2) choosing suitable data sources, (3) setting fair and realistic expectations when recruiting annotators, and (4) avoiding cultural misrepresentation.
@@ 2 Related Work
Work on ethical practices in AI, ML, and NLP research spans a wide range of topics, including artefact documentation Bender (2011); Bender and Friedman (2018); Gebru et al. (2021); Rogers et al. (2021); Mohammad (2022) and best practice recommendations Hollenstein et al. (2020); Mohammad (2023). Research specifically focused on low-resource languages tends to address the general state of NLP in this area Held et al. (2023); Joshi et al. (2020); Blasi et al. (2022); Doğruöz and Sitaram (2022), data collection challenges Yu et al. (2022), limitations in specific tasks such as machine translation Mager et al. (2023), developments in LLM research Mihalcea et al. (2024), and the fundamental issue of including the people whose languages are being studied Mager et al. (2023); Bird (2020, 2022); Bird and Yibarbuk (2024); Lent et al. (2022). Such work highlights the peculiarities of many low-resource languages, the majority of which are vernacular rather than institutionalised or written Bird and Yibarbuk (2024); Bird (2024). It further advocates for language communities to assume agency over their own languages Schwartz (2022); Markl et al. (2024); Mihalcea et al. (2024). For instance, Bird and Yibarbuk (2024) examine how experts, such as linguists and computer scientists, collaborate with language communities through participatory design approaches Winschiers-Theophilus et al. (2010), while Cooper et al. (2024) offer guidance on engaging with Indigenous communities beyond concerns of mere accuracy. Doğruöz and Sitaram (2022) shed light on the need to avoid treating NLP for low-resource languages as a scaled-down version of high-resource language technologies, emphasising the importance of accounting for linguistic and cultural peculiarities. Similarly, Adebara and Abdul-Mageed (2022) stress the significance of such language-specific features (e.g., tones) with a focus on African languages. Beyond language speakers, other work considers the needs of users more broadly. For instance, Blaschke et al. (2024) address the concerns of dialect speakers and emphasise the importance of involving them in the development of language tools and resources. In addition, Yang et al. (2024) define the concept of social awareness and advocate against treating language purely as a computational problem in NLP.
In this paper, we contribute to this ongoing discussion by shifting focus to the practical challenges faced by NLP researchers and practitioners working on mid- and low-resource languages by drawing on methods from social sciences Cetina (1999). We investigate the methodological practices and issues currently shaping the field. To the best of our knowledge, little research has examined how NLP work on low-resource languages engages with online communities, apart from a few case studies involving participatory frameworks (e.g., Masakhane) Birhane et al. (2022), and the work of Lent et al. (2022), who analyse 38 responses from Creole speakers about their experiences with language technologies. Based on an analysis of feedback from our survey respondents, we offer practical recommendations that prioritise transparency and ethically grounded approaches to building more human-centered NLP artefacts for under-served languages.
@@ 3 Survey
Our main goal was to investigate the current challenges and problematic practices in NLP research for mid- to low-resource languages and to propose potential solutions. To this end, we reached out to the NLP community (i.e., *CL networks) between June and October 2024 via platforms such as X (formerly Twitter), LinkedIn, Google Groups, Slack channels of online NLP communities, and direct emails. We specifically targeted researchers and annotators working on mid- and low-resource languages, language variants, dialects, and vernaculars, to survey how research in these areas is conducted. Participants reported on common practices, motivations, and key issues. We then conducted both quantitative and qualitative analyses of the responses.
3.1 Respondents
The respondents are NLP researchers and practitioners involved in data collection, annotation, model development, or other research questions related to under-served languages.
3.2 Survey Structure
We ask the respondents about (1) their previous experiences in the area, (2) current problems and limitations related to their language(s) of interest, (3) the motivation behind their involvement in various projects, and (4) how they were credited for tasks often specific to low-resource language research—such as annotation work conducted via online community forums or participatory frameworks.
Note that we allowed respondents to determine for themselves what constitutes a low-resource language, as there is no universally accepted definition. For example, most researchers would agree that Tamasheq is a low-resource language, whereas opinions may differ regarding Malaysian.
3.2.1 General Questions
Respondents had the option to provide their names and contact information for potential follow-up. They were asked about:
  • the language(s) they work on,
  • the project(s) they have been involved in,
  • whether they are or were part of any online community (i.e., participatory research framework),
  • whether the project(s) they worked on were based in industry, academia, or both,
  • the kinds of NLP tools that are or would be relevant and useful for their language(s) of interest,
  • their reasons for working on this/these language(s).
3.2.2 Reporting on Incentives and Potential Limitations
We looked into the common reasons why researchers work on low-resource languages. Therefore, we asked the participants to report on:
  • the incentive(s) for working on their language(s) of interest,
  • the incentive(s) for working on specific project(s) or task(s).
As we are aware of potential drawbacks in NLP for mid- to low-resource languages Blasi et al. (2022), we examined whether the respondents had been working in the area due to any limitations observed in available NLP tools in their language(s) of interest. Note that these questions were optional as researchers may work on any language for various other reasons. We asked the participants to report on:
3.2.3 Reporting on Credit Attribution
We asked the respondents about how often they were properly credited for their work. Moreover, since reaching out to online participatory frameworks is common to projects that focus on under-resourced languages, we asked whether the participants were involved in past projects through such frameworks. This is because involving communities in NLP and ML projects is relatively new to the field and can therefore be abused (purposefully or not) due to the lack of clear standards regarding data workers in such contexts. Therefore, our questions were the following:
4 Findings
We received 81 responses from researchers working on a wide range of languages and language families. Even though including contact information was optional, more than 90% of the respondents chose not to reply anonymously, and 80% asked for updates on the project. Table 1 shows the distribution of responses to questions on project affiliations, the tasks in which the respondents were involved, and their motivations for working on under-resourced languages. Note that percentages do not sum up to 100% as respondents could report on more than one project. That is, participants could also be involved in several languages and NLP tasks. As shown in Table 1, most participants were involved in dataset curation mainly motivated by scientific interest or curiosity, and for building language technologies because of observed limitations in resources dedicated to their language(s) of interest.
@@ 4.1.1 Projects
Respondents could report on one or multiple projects they had been involved in. As shown in Table 1, most respondents reported working on academic projects. Around one-third participated in collaborations between industry and academia, or were involved in both types of projects. [^1] Note that although over 50% of respondents named the projects they participated in, we do not disclose these in order to protect their anonymity.
@@ 4.1.2 Languages
Among the 81 responses, respondents reported working on over 70 low-resource languages, which they specifically named (see Appendix). Figure 2 illustrates the main regions where these languages are spoken. These include variants, dialects, and vernaculars (e.g., country-specific Arabic dialects), mid- to low-resource languages (e.g., Indonesian), as well as widely acknowledged low-resource languages such as Welsh, Yoreme Nokki, and Setswana. Additionally, around 12% of respondents reported working on language families or branches, such as South Asian languages, all Gaeilige dialects, or Arabic/English varieties. A significant proportion of respondents also work on high-resource languages in parallel.
@@ 4.2 Incentives and Potential Limitations
When asking respondents why they had worked on NLP for under-resourced languages, we provided a checklist from which they could select multiple options and add their own entry. We report on common motivations and practices that are typically specific to mid- to low-resource settings, often due to factors such as data scarcity. We also identify problematic instances and analyse the potential reasons behind some of them.
@@
When further examining the common motivations, we report more detailed numbers in Figure 3. Among those who were motivated by scientific curiosity or interest in Table 1 there were those whose interest was in NLP/CL/ML research (68%) and those whose interest was in languages (68%). Note that the two are not mutually exclusive. For the respondents whose motivation was to build language technologies, most of them were more interested in building technologies for their own language(s) (60%) as opposed to building technologies for as many languages as possible (38%). This is particularly interesting as it constitutes evidence of the power of language as a symbolic capital Bourdieu (1991), which can sometimes manifest in the feeling of “a duty” that one has towards their language.
Other frequent motivations include marked limitations in language resources and tools in the language(s) of interest (60%) and the willingness to contribute to LLM research (59%).
@@ 4.2.2 Qualitative Analysis of the Limitations
We provided the respondents with free text sections where they could report examples of tools or resources that suffer from the limitations they mentioned to justify their choices. When manually processing the answers, we noticed the following recurring topics:
@@
    Limitations related to existing resources: such as their public unavailability, small size, or limited representativeness and quality.
@@
    Fundamental problems related to NLP research on mid- to low-resource languages: such as the lack of funding often due to the “low prestige” language dilemma—the false notion that some languages or language varieties are more important than others. Interestingly, this was equally observed in projects from academia and industry.
@@
Lack of representativeness and unnaturalness of the data were commonly reported in the responses. The respondents reported a lack of awareness of language variants and cultural aspects when building a language-specific artefact; the reliance on the standardised version of a given language due to power dynamics (i.e., more power in the hands of well-funded institutions and established researchers); the presence of offensive utterances in the data due to a lack of data filtering; and potentially wrong assumptions about a language or a culture. For instance, the time-specific context and usage of some languages, such as ancestral ones (e.g., Coptic), have considerably changed and one has to take these facts into account. In addition, datasets may be collected from inadequate sources or could be aligned with Western values, standards, or expectations. This can be due to power differentials or a lack of deeper examination carried along with locals and native speakers. Finally, researchers rely on personal connections as it is hard to impossible to find fluent speakers of mid- to low-resource languages on commonly used annotation platforms such as Amazon Mechanical Turk and Polific. Added to this reason, the lack of funding leads researchers to turn to participatory frameworks. This practice has been at the center of major NLP contributions in recent years Birhane et al. (2022). However, despite its benefits for people with common research interests, the absence of well-established standards puts vulnerable community members at risk as their efforts may not be properly recognised.
@@
Many respondents reported that conducting research in mid- to low-resource languages often entails high costs of data curation and potential outreach to local communities. Further, when resources for an under-served language exist, they are often not freely available.